To what value can RE be reduced without transistor going into saturation. Given : Vcc = 5V, Rc = 1kΩ, RB = 22kΩ, RE = 2.2kΩ, -VEE = -5V.
1
Expert's answer
2020-10-05T10:54:58-0400

Let "\beta = 100" and "V_{RE} = 3.9 V"

"I_E = \frac{V_{EE}}{R_E + \frac{R_B}{\beta}} = 2.06 mA"

"V_C = V_{CC} - I_CR_C = 2.94 V"


At saturation, "V_{C(sat)} = V_{E(sat)}"

"I_{C(sat)} = \frac{V_{CC}-V_{C(sat)}}{R_C} = 10 mA"


"(R_{E})_{min} = \frac{V_{RE}}{I_{E(sat)}} = 390 \Omega"


"\Delta R_E = 2200 - 390 = 1810\Omega"